# The State of Open Science in Software Engineering Research: A Case Study of ICSE Artifacts

### v2

@@ 4.4. Answering RQ4: Reproducibility Outcomes and Contributing Factors
Figure 7 shows the distribution of reproducibility outcomes. While RQ1 revealed executability challenges, reproducibility results are even more concerning. Only 8% of artifacts were fully reproducible without modification, 5% with minor adjustments, and 1% with major effort, yielding a total reproducibility rate of just 14%. Another 16% were partially reproducible, often requiring major or critical modifications (62.5%). These results emphasize that successful execution does not ensure reproducibility. Fully reproducible artifacts were better documented and required lower effort (85.71% low to moderate), while 50% of partially reproducible ones required high effort.
Among executable artifacts, only 35% reproduced results matching the original study. Another 30% produced divergent outputs, and 5% produced no results at all. The situation was worse for partially executable artifacts, where 45.83% failed to generate any usable results due to incomplete pipelines, missing components, runtime crashes, or logical errors.
A major barrier was the absence of validation guidance. For 30% of executable artifacts and 20.83% of partially executable ones, we could not verify results despite execution, due to missing instructions or unclear output structures. Only 6.25% provided sufficient validation steps for comparison.
@@ 5.1. Actionable Guidelines for Improving Executability
G1: Provide Comprehensive and Structured Documentation. Findings from RQ2 and RQ3 reveal that well-structured documentation is one of the key determinant of artifact executability and reproducibility. Our analysis identified five essential documentation components:
@@
G3: Combine Containerized Environments with Accessible Source Code. Containerization encapsulates the development environment and dependencies, improving portability and reducing setup complexity, while source code ensures transparency, flexibility, and long-term adaptability. In our dataset, 34 artifacts included both containerized environments (e.g., Docker, VM) and raw source code. Artifacts with containerized setups achieved higher executability (67.47%) and reproducibility (52.17%) compared to source-only artifacts (31% and 33.25%, respectively). Most containerized artifacts (83.33%) ran successfully with low to moderate effort, and 75% required no modification, indicating that encapsulated environments substantially simplify execution. This becomes particularly critical when executing a study that requires complex setup procedures and is highly sensitive to environmental variations. However, containerization introduces trade-offs. When execution failed, 57.14% of containerized artifacts were unrecoverable due to restricted access or opaque configurations. In contrast, raw source code, though more demanding to configure, offers greater transparency for debugging, reuse, and extension.
@@ 5.2. Preparation and Evaluation of Our Replication Package
Preparation of our Replication Package. Our replication package was constructed following the actionable guidelines derived from executing and reproducing 100 ICSE replication packages in this study. It is organized into two complementary parts. Part 1 contains all analysis scripts and datasets required to regenerate the results reported in this paper. It was developed in full compliance with the documentation and packaging guidelines introduced in Section 5.1. Along with the source code, we also provide a Dockerfile to run our experiments in an isolated environment. Part 2 provides a curated, per-artifact record of our execution and reproduction attempts for 100 ICSE replication packages. It does not modify the original artifacts but documents our interactions with them to ensure full traceability and transparency. Each artifact record includes four files (README.md, execution-log.md, patch.md, and updated-instructions.md) and one folder (/errors) when applicable. The detail purpose of these files can be found in the README.md file of our replication package (Replication Package, 2026).
@@
$\bullet$ Assessment Summary (Part One). All four evaluators executed part one in full using both the Jupyter notebooks and the Docker setup, following the provided instructions. They evaluated documentation completeness, executability, reproducibility, and overall usability in accordance with our proposed guidelines. The evaluators unanimously confirmed that the package contained all five essential documentation components (Project Metadata, System Specification, Installation & Environment Setup, Execution Steps, and Validation Steps), thereby fully satisfying G1 (comprehensive documentation). For G2 (avoiding configuration leakage), all evaluators successfully reproduced the results without encountering any modifications or environment-related issues. To further strengthen this assessment, the first author independently validated Part One in clean, isolated environments across multiple operating systems, confirming the absence of hidden dependencies or environment leakage.
@@
$\bullet$ Assessment Summary (Part Two). Table 3 summarizes the assessment of Part II by the four evaluators (individual responses are available in our replication package (Replication Package, 2026)). We evaluated this part using a stratified random sample of 20 artifacts drawn proportionally from the three executability categories (8 Executable, 5 Partially Executable, and 7 Not Executable) and distributed evenly among four evaluators. Each evaluator first reviewed the README.md file to verify the factual accuracy of our documentation. They then executed each artifact twice: once strictly following the original instructions and once using our meta-documentation (execution-log.md, patch.md, and updated-instructions.md). This two-pass execution allowed evaluators to assess both the baseline usability of the artifacts and the effectiveness of our structured documentation in improving comprehensibility and execution success.
@@
All evaluators reproduced our reported results, and the final executability and reproducibility status remained unchanged across all 20 artifacts. Nineteen artifacts were fully consistent with our recorded classifications. The single disagreement occurred because one evaluator’s local machine already contained a dependency (build-essential) that our clean Docker setup required to be installed explicitly. Evaluators rated 18 artifacts as factually accurate and 2 as partially accurate, noting only minor corrections that were later re-checked by the authors. In 12 cases, our meta-documentation was judged to be clearer and more usable than the original instructions. Usability remained unchanged for 5 artifacts whose original documentation was already sufficient. The remaining 3 artifacts were not executable, and their replication packages contained no execution instructions, making it impossible for us to improve their execution processes further.
Although we could not directly apply our proposed guidelines (Section 5.1) to third-party artifacts, we compensated through structured meta-documentation that improved traceability and reproducibility. In line with G1, we evaluated each artifact against the five documentation components and filled missing details in our meta-files, which evaluators later confirmed as factually accurate. For G2, since we could not control environment leakage in external packages, we mitigated it by recording complete execution logs, patches, and errors and providing them in our replication package in a structured, hyperlinked format. Regarding G3, re-containerizing all 100 heterogeneous artifacts was infeasible; instead, we documented key environmental and configuration details, which evaluators reported as improving clarity and reusability except in cases where the original packages lacked sufficient information for full replication.
Given that our replication package consists of two interrelated parts, evaluating the main README.md file was essential to ensure users could understand the study design, the role of each part, and the steps to reuse our artifact. Evaluators assessed it for documentation quality, linkage among components, and ease of navigation, awarding a perfect average score of 5/5 for clarity and completeness. Minor suggestions were carefully incorporated into the final version of our replication package. Overall, the combination of industry and academic evaluators provided balanced evidence that our replication package is usable in practice, transparent in structure, and methodologically sound for research reuse.
@@ 5.3. Implications
Our study provides actionable insights for multiple stakeholders in the SE community, including researchers, conference organizers, and industry practitioners.
For Researchers, the findings of this study can help develop more usable and reproducible replication packages. Researchers can follow our proposed documentation structure to ensure clear execution and validation paths, and they can provide both source code and containerized environments to balance usability and extensibility. They can also apply practical strategies derived from our experience, such as using version-buffer heuristics when version details are missing. For Conference Organizers, our results can inform improved artifact submission and evaluation policies. For example, asking reviewers to validate packages in isolated environments (such as fresh Docker containers) can reveal hidden configuration issues and better reflect real-world usability. For the SE Community, the shortcomings we identified highlight the need for dedicated platforms for hosting and maintaining research artifacts, ideally with automated quality checks that detect common executability and reproducibility barriers and guide authors in following best practices to advance open science in SE reseach.
@@ 6. Threats to Validity
Threats to external validity refer to the extent to which our findings generalize beyond the studied sample. Our dataset comprises replication packages from ICSE research track papers, which may limit generalizability to other SE venues or domains. However, ICSE is one of the most prestigious SE conferences, with rigorous artifact evaluation processes, diverse SE topics, and consistently high-quality publications. As a result, our findings likely represent a best-case scenario for open science practices in the field. This study can be extended in the future to other major venues such as FSE and ASE to examine whether similar patterns appear across the broader SE research community. A second external validity threat concerns temporal generalizability, as we examine a 10-year window (2015–2024). This period is broad enough to capture evolving artifact practices but may not reflect earlier trends. Expanding the window further would dilute yearly samples and reduce representativeness. To mitigate this, we applied stratified random sampling across years to ensure balanced coverage.
Threats to internal validity relate to potential errors in data collection and analysis. Our evaluation involved manual steps, including artifact identification, executability assessment, effort classification, and reproducibility evaluation. We mitigated human error through a multi-pass review process, and ambiguous cases were independently reviewed by multiple authors to reach consensus. Evaluators’ varying technical expertise could also introduce bias, especially when interpreting ambiguous instructions. To address this, evaluators prioritized strict adherence to the original documentation, using domain expertise only when execution failed under the provided guidance.
@@
Threats to construct validity concern the correctness and consistency of how we measured executability, effort, and reproducibility. Although these constructs are grounded in prior literature, some steps, such as identifying modification types or diagnosing failure causes, require judgment. We mitigated subjectivity bias through structured coding schemes, predefined criteria, and agreement across evaluators. When one evaluator could not execute an artifact, a second evaluator repeated the process independently. If both encountered the same issue, we attributed the failure to the artifact. Our effort estimates, based on agile-style bucketed time ranges, may be sensitive to changes in the boundaries. A sensitivity analysis showed that increasing the ranges by 5-10 percent produced no changes, while decreasing them caused only minor shifts, such as a few low-effort cases moving to moderate, indicating a negligible impact on overall findings. Hardware constraints present another construct validity threat. Some artifacts required extremely high-end setups, such as machines with 250 GB of RAM or multiple NVIDIA A100 GPUs, which we could not precisely replicate. We mitigated this limitation by using the closest available hardware (Section 3.5) and adjusting permitted configurations, such as batch size and dependency versions, accordingly.
@@ 7. Related Work
Open science has gained increasing traction in SE, with many studies proposing frameworks, tools, and guidelines to address reproducibility and replicability concerns. This section synthesizes the most relevant work across four core themes aligned with our investigation.
Artifact Sharing and Availability. Studies emphasize that open artifact sharing is essential. Oliveira Jr. et al. (Oliveira Jr et al., 2024) and Cordeiro and Oliveira Jr. (Cordeiro and Oliveira Jr, 2025a, b) advocate SE-specific frameworks and show that artifact badges improve availability. Liu et al. (Liu et al., 2024) confirm a rising trend in sharing but warn that availability alone does not ensure usability. Others highlight reproducible research as a standard practice (Madeyski and Kitchenham, 2017) and call for structured evaluations, especially in MSR (Tkachenko, 2025; Gonzalez-Barahona and Robles, 2023).
Executability Challenges. Despite wider sharing, executability remains difficult. Mukhtar et al. (Mukhtar et al., 2024) found missing dependencies and poor documentation in SE tools. Ajayi et al. (Ajayi et al., 2023) and Bonneel et al. (Bonneel et al., 2020) report failures due to configuration drift. Proposed solutions include containerization (Costa et al., 2024) and modeling deep environment variability (Acher et al., 2024).
Reproducibility Audits. Several audits show that execution does not imply reproducibility. Nong et al. (Nong et al., 2023), Agal et al. (Agal et al., 2025), and Liu et al. (Liu et al., 2022) cite inconsistent pipelines and missing criteria. Earlier, Mahmood et al. (Mahmood et al., 2018) and Rodriguez-Perez et al. (Rodríguez-Pérez et al., 2018a) highlighted gaps in defect prediction reproducibility. Robles and Gonzalez-Barahona (Gonzalez-Barahona and Robles, 2023) proposed structured MSR assessment protocols.
Community Standards and Taxonomies. Initiatives like the ACM artifact badge clarify reproducibility terms. Gundersen and Kjensmo (Gundersen and Kjensmo, 2018a) introduced a taxonomy; Cordeiro and Oliveira Jr. (Cordeiro and Oliveira Jr, 2025a) emphasized usability and clarity. Acher et al. (Acher et al., 2024) stress modeling software variability for reproducibility.
@@ 8. Conclusion
Despite the growing emphasis on open science in SE research, the practical usability of shared artifacts remains insufficiently examined. This study provides the first large-scale empirical assessment of executability and reproducibility across 100 ICSE replication packages. By investing roughly 650 person-hours of analysis, we uncovered substantial gaps; only 40% of artifacts were executable, just one-third ran without modification, and reproducibility was achieved in only 35% of those that executed. To help close this gap, we proposed three actionable guidelines; adopt comprehensive documentation, validate artifacts in clean environments to avoid configuration leakage, and supply both containerized setups and accessible source code (where applicable). Our findings show that artifacts including these components achieve substantially higher executability and reproducibility. Strengthening open science in SE requires not only sharing artifacts but ensuring they are runnable, verifiable, and usable by others. Following the guidelines introduced in this study can help researchers and conference organizers advance reproducible and trustworthy SE research.

### v4

@@ 4.4. Answering RQ4: Reproducibility Outcomes and Contributing Factors
Figure 7 shows the distribution of reproducibility outcomes. While RQ1 revealed challenges in executability, reproducibility results are even more concerning. Only 8% of artifacts were fully reproducible without modification, 5% with minor adjustments, and 1% with major effort, yielding a total reproducibility rate of just 14%. Another 16% were partially reproducible, often requiring major or critical modifications (62.5%). These results emphasize that successful execution does not ensure reproducibility. Fully reproducible artifacts were better documented and required less effort (85.71% low to moderate), whereas 50% of partially reproducible artifacts required high effort.
Among the 40 fully executable artifacts, only 35% reproduced results matching the original study. Another 30% produced divergent outputs, and 5% produced no results. The situation was worse for partially executable artifacts, where 45.83% failed to generate any usable results due to incomplete pipelines, missing components, runtime crashes, or logical errors. A major barrier was the absence of validation guidance. For 30% of executable artifacts and 20.8% of partially executable artifacts, we could not verify the results despite execution due to missing instructions or unclear output structures. Only 6.3% provided sufficient validation steps for comparison.
@@ 5.1. Actionable Guidelines for Improving Executability
G1: Provide Comprehensive and Structured Documentation. Findings from RQ2 and RQ3 reveal that well-structured documentation is one of the key determinants of artifact executability and reproducibility. Our analysis identified five essential documentation components:
@@
G3: Combine Containerized Environments with Accessible Source Code. Containerization encapsulates the development environment and dependencies, improving portability and reducing setup complexity, while source code ensures transparency, flexibility, and long-term adaptability. In our dataset, 34 artifacts included both containerized environments (e.g., Docker, VM) and raw source code. Artifacts from containerized setups achieved higher executability (67.47%) and reproducibility (52.17%) than source-only artifacts (31% and 33.25%, respectively). Most containerized artifacts (83.33%) ran successfully with low to moderate effort, and 75% required no modification, indicating that encapsulated environments substantially simplify execution. This becomes particularly critical when executing a study that requires complex setup procedures and is highly sensitive to environmental variations. However, containerization introduces trade-offs. When execution failed, 57.14% of containerized artifacts were unrecoverable due to restricted access or opaque configurations. In contrast, raw source code, though more demanding to configure, offers greater transparency for debugging, reuse, and extension.
@@ 5.2. Preparation and Evaluation of Our Replication Package
Preparation of our Replication Package. Our replication package was constructed following the actionable guidelines derived from executing and reproducing 100 ICSE replication packages in this study. It is organized into two complementary parts. Part 1 contains all analysis scripts and datasets required to regenerate the results reported in this paper. It was developed in full compliance with the documentation and packaging guidelines introduced in Section 5.1. Along with the source code, we also provide a Dockerfile to run our experiments in an isolated environment. Part 2 provides a curated, per-artifact record of our execution and reproduction attempts for 100 ICSE replication packages. It does not modify the original artifacts but documents our interactions with them to ensure full traceability and transparency. Each artifact record includes four files (README.md, execution-log.md, patch.md, and updated-instructions.md) and one folder (/errors) when applicable. The detailed purpose of these files can be found in the README.md file of our replication package (Muttakin et al., 2026).
@@
$\bullet$ Assessment Summary (Part One). All four evaluators executed part one in full using both the Jupyter notebooks and the Docker setup, following the provided instructions. They evaluated documentation completeness, executability, reproducibility, and overall usability in accordance with our proposed guidelines. The evaluators unanimously confirmed that the package contained all five essential documentation components (Project Metadata, System Specification, Installation & Environment Setup, Execution Steps, and Validation Steps), thereby fully satisfying G1 (comprehensive documentation). For G2 (avoiding configuration leakage), all evaluators successfully reproduced the results without any modifications or environment-related issues. To further strengthen this assessment, the first author independently validated Part One in clean, isolated environments across multiple operating systems, confirming the absence of hidden dependencies or environment leakage.
@@
$\bullet$ Assessment Summary (Part Two). Table 3 summarizes the assessment of Part II by the four evaluators (individual responses are available in our replication package (Muttakin et al., 2026)). We evaluated this part using a stratified random sample of 20 artifacts drawn proportionally from the three executability categories (8 Executable, 5 Partially Executable, and 7 Not Executable) and distributed evenly among four evaluators. Each evaluator first reviewed the README.md file to verify the factual accuracy of our documentation. They then executed each artifact twice: once strictly following the original instructions and once using our meta-documentation (execution-log.md, patch.md, and updated-instructions.md). This two-pass execution allowed evaluators to assess both the baseline usability of the artifacts and the effectiveness of our structured documentation in improving comprehensibility and execution success.
@@
All evaluators reproduced our reported results, and the final status of executability and reproducibility remained unchanged across all 20 artifacts. Nineteen artifacts were fully consistent with our recorded classifications. The single disagreement occurred because one evaluator’s local machine already contained a dependency (build-essential) that our clean Docker setup required to be installed explicitly. Evaluators rated 18 artifacts as factually accurate and 2 as partially accurate, noting only minor corrections that were later re-checked by the authors. In 12 cases, our meta-documentation was judged to be clearer and more usable than the original instructions. Usability remained unchanged for 5 artifacts whose original documentation was already sufficient. The remaining 3 artifacts were not executable, and their replication packages contained no execution instructions, so we could not further improve their execution processes.
Although we could not directly apply our proposed guidelines (Section 5.1) to third-party artifacts, we compensated by providing structured meta-documentation, which improved traceability and reproducibility. In line with G1, we evaluated each artifact against the five documentation components and filled missing details in our meta-files, which evaluators later confirmed as factually accurate. For G2, since we could not control environment leakage in external packages, we mitigated it by recording complete execution logs, patches, and errors and providing them in our replication package in a structured, hyperlinked format. Regarding G3, re-containerizing all 100 heterogeneous artifacts was infeasible. Instead, we documented key environmental and configuration details that evaluators reported as helpful for improving clarity and reusability, except in cases where the original packages lacked sufficient information for full replication.
Given that our replication package consists of two interrelated parts, evaluating the main README.md file was essential to ensure that users could understand the study design, the roles of each part, and the steps for reusing our artifact. Evaluators assessed it for documentation quality, inter-component linkage, and ease of navigation, awarding a perfect average score of 5/5 for clarity and completeness. Minor suggestions were carefully incorporated into the final version of our replication package. Overall, the combination of industry and academic evaluators provided balanced evidence that our replication package is usable in practice, transparent in structure, and methodologically sound for research reuse.
@@ 5.3. Implications
Our study provides actionable insights for multiple stakeholders in the SE community, including researchers, conference organizers, and industry practitioners. For Researchers, the findings of this study can help develop more usable and reproducible replication packages. Researchers can follow our proposed documentation structure to ensure clear execution and validation paths, and they can provide both source code and containerized environments to balance usability and extensibility. They can also apply practical strategies derived from our experience, such as using version-buffer heuristics when version details are missing. For Conference Organizers, our results can inform the development of improved artifact submission and evaluation policies. For example, asking reviewers to validate packages in isolated environments (e.g., fresh Docker containers) can reveal hidden configuration issues and more accurately reflect real-world usability. For the SE Community, the shortcomings we identified underscore the need for dedicated platforms to host and maintain research artifacts, ideally with automated quality checks that detect common barriers to executability and reproducibility and guide authors in following best practices to advance open science in SE research.
@@ 6. Threats to Validity
Threats to external validity refer to the extent to which our findings generalize beyond the studied sample. Our dataset comprises replication packages from ICSE research-track papers, which may limit its generalizability to other SE venues or domains. However, ICSE is one of the most prestigious SE conferences, with rigorous artifact evaluation processes, a diverse range of SE topics, and consistently high-quality publications. As a result, our findings likely represent a best-case scenario for open science practices in the field. This study can be extended to other major venues, such as FSE and ASE, to examine whether similar patterns appear across the broader SE research community. A second external validity threat concerns temporal generalizability, as we examine a 10-year window (2015-2024). This period is broad enough to capture evolving artifact practices but may not reflect earlier trends. Expanding the window further would dilute yearly samples and reduce representativeness. To mitigate this, we applied stratified random sampling across years to ensure balanced coverage.
Threats to internal validity relate to potential errors in data collection and analysis. Our evaluation involved several manual steps, including artifact identification, executability assessment, effort classification, and reproducibility evaluation. To reduce human error, we employed a multi-pass review process, and ambiguous cases were independently examined by multiple authors to reach consensus. Evaluator expertise may influence executability outcomes. Although two experienced authors independently attempted execution while strictly adhering to the original documentation, differences in technical background, tool familiarity, or system configurations may lead other practitioners to obtain different results. Thus, artifacts classified as non-executable should be interpreted as unsuccessful under controlled and informed conditions rather than as evidence of absolute infeasibility.
@@
Threats to construct validity concern the correctness and consistency of how we measured executability, effort, and reproducibility. While these constructs are commonly used in empirical studies, some steps, such as identifying modification types or diagnosing failure causes, require judgment. We mitigated subjectivity bias through structured coding schemes, predefined criteria, and agreement across evaluators. When one evaluator could not execute an artifact, a second evaluator repeated the process independently. If both encountered the same issue, we attributed the failure to the artifact. Our effort estimates, based on agile-style bucketed time ranges, may be sensitive to changes in the boundaries. A sensitivity analysis showed that increasing the ranges by 5-10 percent produced no changes, while decreasing them caused only minor shifts, such as a few low-effort cases moving to moderate, indicating a negligible impact on overall findings. Hardware constraints present another construct validity threat. Some artifacts required extremely high-end setups, such as machines with 250 GB of RAM or multiple NVIDIA A100 GPUs, which we could not precisely replicate. We mitigated this limitation by using the closest available hardware (Section 3.5) and adjusting permitted configurations, such as batch size and dependency versions, accordingly.
@@ 7. Related Work
Open science has gained increasing traction in SE, with several studies promoting it as a foundation for rigorous research, empirical investigations, and technical and methodological solutions to improve research transparency and reproducibility. Oliveira Jr et al. (2024) proposed a conceptual framework that treats open science as a requirement for promoting a cultural shift within the SE community. Cordeiro and Oliveira Jr (2025b) examined how reproducibility is defined, evaluated, and supported, while others highlight reproducible research as a scientific standard for addressing validity concerns in SE (Madeyski and Kitchenham, 2017; Tkachenko, 2024).
Prior studies have also assessed the reproducibility of SE research to identify existing practices, challenges, and gaps. Liu et al. (2024) report an upward trend in artifact sharing but caution that availability alone does not guarantee usability. Similarly, Cordeiro and Oliveira Jr (2025a), based on a survey of SE researchers, found that limited data sharing and incomplete experimental details remain key barriers to reproducibility. Beyond these issues, several technical factors may further hinder reproducibility, including dependency problems, incomplete documentation, and insufficient information to reproduce results (Ajayi et al., 2023; Bonneel et al., 2020; Rodríguez-Pérez et al., 2018). In deep learning-based SE studies, additional challenges arise from inconsistent evaluation practices, unstable experimental setups, and missing details about baselines, preprocessing steps, and hyperparameters (Nong et al., 2022; Liu et al., 2021; Mukhtar et al., 2024).
Researchers have also explored technical and methodological solutions to improve the availability and reproducibility of SE research. Méndez Fernández et al. (2019) introduced an open science initiative for the Empirical Software Engineering journal, while Gonzalez-Barahona and Robles (2023) proposed structured assessment protocols for Mining Software Repository studies. Mahmood et al. (2018) further recommended improved reporting guidelines and incentive mechanisms to strengthen replication efforts. Additional technical solutions include containerization to lock execution environments (Costa et al., 2026) and modeling environment variability to identify potential sources of irreproducibility (Acher et al., 2024).
@@ 8. Conclusion
Despite the growing emphasis on open science in SE research, the practical usability of shared artifacts has received insufficient attention. In this study, we provide a comprehensive empirical assessment of executability and reproducibility across 100 ICSE replication packages. By investing roughly 650 person-hours of analysis, we uncovered substantial gaps. Only 40% of artifacts were executable, about one-third of these ran without modification, and reproducibility was achieved in only 35% of the executable artifacts. To help close this gap, we proposed three actionable guidelines: adopt comprehensive documentation, validate artifacts in clean environments to avoid configuration leakage, and supply both containerized setups and accessible source code (where applicable). Our findings show that artifacts that include these components achieve substantially higher levels of executability and reproducibility. Strengthening open science in SE requires not only sharing artifacts but ensuring they are runnable, verifiable, and usable by others. Following the guidelines introduced in this study can help researchers and conference organizers advance reproducible and trustworthy SE research.